No friction on horizontal surface

the two plates are the same (drawing). Symbols of lines and

angular velocity as the side, dimensions R and r (R> r), va

Touch is elastic. Determine the rear disc speed

touch.


1
Expert's answer
2021-02-25T17:53:27-0500

We have given radius of two plates as r and R


According to Impulse momentum theorem,


Impulse = Change in momentum


P =mv, where v is the linear velocity


Using relation between angular and linear velocity  "v=\omega\times r" ,with the angular velocity


"v=\omega r" , where "\omega" is the angular velocity P="m \omega R"


P="m \omega R"  as r=R


r=R in above problem therefore "w=\dfrac{P}{mR}"


"w=\dfrac{P}{mR}"  is the angular velocity of plate of radius R